Bishop, Inc., is obligated to pay its creditors $6.200 during the year. (Leave no cells blank - be certain to enter O wherever req ed.) a. What is the market value of the shareholders equity if assets have a market value of $8,300? (Do not round intermediate calculations.) b. What is the market value of the shareholders equity if assets equal $5,600? (Do not round intermediate calculations.) ook int a. Market value b. Market value ences
Answer #1

Total assets=Total equity+Total liabilities

1.Shareholders equity=(8300-6200)=$2100

2.Shareholders equity=(5600-6200)=0